### v1

# Finite-sample valid prediction of future insurance claims in the regression problem

Abstract
In the current insurance literature, prediction of insurance claims in the regression problem is often performed with a statistical model. This model-based approach may suffer from several drawbacks: (i) model misspecification, (ii) selection effect, and (iii) lack of finite-sample validity. This article addresses these three issues simultaneously by employing conformal prediction—a general machine learning strategy for valid predictions. The proposed method is both model-free and tuning-parameter-free. It also guarantees finite-sample validity at a pre-assigned coverage probability level.
Keywords and phrases: conformal prediction; explainable machine learning; insurance data science; predictive analytics.
1 Introduction
Prediction is one of the most important inferential tasks for actuaries since it forms the basis for many key aspects of an insurer’s business operations, such as premium calculation and reserves estimation. According to Shmueli (2010), there are two key goals in data science and statistics: to explain and to predict. However, these two goals often warrant different approaches. For example, as demonstrated in Shmueli (2010), a wrong model, under some conditions, can even beat the oracle model in prediction, but the same cannot be said for explanation. This paper only concerns prediction. In the existing insurance literature, prediction is often performed using either a parametric approach or a non-parametric approach (e.g., Frees et al. 2014). In the parametric approach, the actuary posits a model, applies model selection tools to choose the “best” model, trains the chosen model, and finally makes predictions; see, for example, Claeskens and Hjort (2008) and Part I of Frees (2010). While this parametric approach has been widely applied in insurance, it has several drawbacks. First, the posited model may be misspecified, leading to grossly misleading predictions (Hong and Martin 2020). Second, scholars recently discovered that model selection might introduce serious biases in prediction; this adverse effect is called the selection effect (e.g., Berk et al. 2013; Kuchibhotla et al. 2022); see, for example, Hong and Martin (2018a, b) for some simple illustrations in insurance. Though several proposals have been advanced to alleviate the selection effect, a satisfying answer has yet to be found (Kuchibhotla et al. 2022). Third, prediction is uncertain and should be probabilistically calibrated to achieve finite-sample validity (Gneiting and Katzfuss 2014; Vovk 2024); unfortunately, no parametric model is known to achieve finite-sample validity. For these reasons, some actuaries opt for a non-parametric model for predicting insurance claims (e.g., Jeon and Kim 2013; Hong and Martin 2017; Richardson and Hartman 2018). While most non-parametric models shun both the model misspecification risk and the selection effect, they can still fall short of our expectations for two reasons. In the first place, many of these models contain some tuning parameters and the choices of these tuning parameters are often debatable. In general, there is hardly a choice of the tuning parameter that works across different data sets. If a data-driven method, such as cross-validation, is employed to select tuning parameters, it amounts to throwing the baby out with the bathwater: prediction will be subject to the selection effect. Secondly, predictions made using non-parametric models are usually only asymptotically valid (not finite-sample valid). Since all real-world applications are done in finitely many steps, any asymptotically correct model still leaves much to be desired regarding predictive validity.
Given the above observations, it is natural to ask whether we can develop a predictive method with the following desiderata:
  1. (i)
    it is model-free;
  2. (ii)
    it is tuning-parameter-free;
  3. (iii)
    it provides finite-sample valid predictions.
The answer is affirmative. Indeed, conformal prediction (e.g. Vovk et al. 2005; Shafter and Vovk 2008) is a general machine learning strategy that can be used to create such a predictive method. Hong and Martin (2021) and Hong (2023) introduce conformal predction to insurance for the first time. These two papers develop predictive methods with all the above desiderata for unsupervised learning. The main goal of this paper is to develop an analogous method for the regression problem, i.e., a model-free and tuning-parameter-free method for making finite-sample valid predictions of a continuous response variable in supervised learning.
The remainder of the paper proceeds as follows. Section 2 provides readers with the necessary background by reviewing conformal prediction. Then, Section 3 describes the proposed method in detail. Next, Section 4 uses numerical examples to demonstrate the excellent performance of the proposed method. Finally, Section 5 concludes the paper with concluding remarks. The Appendix contains some technical details.
2 Conformal prediction
Conformal prediction is a general machine learning strategy for generating prediction regions at a pre-determined coverage probability level. Shafer and Vovk (2008) give a user-friendly introduction to conformal prediction. For a comprehensive treatment, see Vovk et al. (2005). Conformal prediction can be used to tackle both supervised learning and unsupervised learning problems. Here we only review conformal prediction in supervised learning.
Suppose we observe pairs $Z_{i}=(X_{i},Y_{i})$, for $i=1,\ldots,n$, where $n$ is the sample size, $X_{i}\in\mathbb{R}^{p}$ is a vector of $p$ features, and $Y_{i}$ is the response variable. Throughout, $p$ is a positive integer. For any $n\geq 1$, we assume the finite sequence $Z_{1},Z_{2},\ldots,Z_{n}$ is exchangeable, i.e., the joint distribution of $(Z_{1},\ldots,Z_{n})$ is the same as the joint distribution of $(Z_{\sigma(1)},\ldots,Z_{\sigma(n)})$, where $\sigma:\{1,2,\ldots,n\}\rightarrow\{1,2,\ldots,n\}$ is any permutation on the set $\{1,2,\ldots,n\}$. Note that the exchangeability assumption is weaker than the widely used iid (independent identically distributed) assumption. In particular, if $Z_{1},\ldots,Z_{n}$ are iid, they are exchangeable (e.g., Chapter 1, Schervish 1995). Throughout, upper-case letters will denote random variables/vectors while the corresponding lower-case letters stand for their realizations.
Our chief goal is to predict the next response $Y_{n+1}$ at a randomly sampled feature $X_{n+1}$, based on past observations of $Z^{n}=\{Z_{1},\ldots,Z_{n}\}$. To apply conformal prediction, the actuary starts with a real-valued deterministic mapping $M$ of two arguments $(B,z)\mapsto M(B,z)$, where the first argument $B$ is a bag, i.e., a collection, of observed data and the second argument $z$ is a provisional value of a future observation to be predicted based on the data in $B$. $M$ is called a non-conformity measure. It measures the degree of non-conformity of the provisional value $z$ with the data in $B$. When the provisional value $z$ agrees with the data in $B$, $M(B,z)$ will be relatively small; otherwise, it will tend to be large. For example, if $z$ is real-valued and $B=\{z_{1},\ldots,z_{n}\}$, then we can take $M(B,z)=|\hat{m}_{B}(x)-y|$, where $\hat{m}_{B}(\cdot)$ is an estimate of the conditional mean function, $\mathsf{E}(Y\mid X=x)$, based on the bag $B$. There is no unique choice of the non-conformity measure. In practice, the non-conformity is chosen at the discretion of the actuary. Once a non-conformity measure is chosen, the actuary runs the conformal prediction algorithm—Algorithm 1 to predict the value of the next label $Y_{n+1}$ at a randomly sampled feature $X_{n+1}=x_{n+1}$.
1 Initialize: data $z^{n}=\{z_{1},\ldots,z_{n}\}$ and $x_{n+1}$, non-conformity measure $M$, and a provisional value $y$ of $Y$;
2 for each possible $y$ do
3       Set $z_{n+1}=(x_{n+1},y)$ and write $z^{n+1}=z^{n}\cup\{z_{n+1}\}$;
4       Define $\mu_{i}=M(z^{n+1}\setminus\{z_{i}\},z_{i})$ for $i=1,\ldots,n,n+1$;
5       Compute $\mathsf{pl}_{z^{n}}(x_{n+1},y)=(n+1)^{-1}\sum_{i=1}^{n+1}1\{\mu_{i}\geq\mu_{n+1}\}$;
6
7 end for
8Return $\mathsf{pl}_{z^{n}}(x_{n+1},y)$ for each possible $y$;
Algorithm 1 Conformal prediction (supervised learning)
In Algorithm 1, $1_{E}$ denotes the indicator function of an event $E$. The quantity $\mu_{i}$, called the $i$-th non-conformity score, assigns a numerical value to $z_{i}$ to indicate how much $z_{i}$ agrees with the data in the $i$-th augmented bag $\widetilde{B}_{i}=z^{n}\cup\{z_{n+1}\}\backslash\{z_{i}\}$, where $z_{i}$ itself is excluded to avoid biases as in leave-one-out cross-validation. The function $\mathsf{pl}_{z^{n}}$ is called the plausibility function. It outputs a value between 00 and $1$, based on all non-conformity scores. The output of $\mathsf{pl}_{z^{n}}$ shows how plausible $z$ is a value of $Z_{n+1}$, based on the available data $Z^{n}=z^{n}$ and $X_{n+1}=x_{n+1}$. Using the output of the plausibility function $\mathsf{pl}_{z^{n}}$, the actuary can construct a $100(1-\alpha)\%$ conformal prediction region
$C_{\alpha}(x;Z^{n})=\{y:\mathsf{pl}_{Z^{n}}(x,y)>\alpha\},$ (1)
where $10<\alpha<$ 1. The next theorem establishes the finite-sample validity of the conformal prediction region in (1).
Theorem 1.
If $\mathsf{P}$ denotes the distribution (i.e., the law) of an exchangeable sequence $Z_{1},Z_{2},\ldots$, then write $\mathsf{P}^{n+1}$ for the corresponding joint distribution of $Z^{n+1}=\{Z_{1},\ldots,Z_{n},Z_{n+1}\}$. For $10<\alpha<$ 1, define $t_{n}(\alpha)=(n+1)^{-1}\lfloor(n+1)\alpha\rfloor$, where $\lfloor a\rfloor$ denotes the greatest integer less than or equal to $a$. Then
$\sup\mathsf{P}^{n+1}\{\mathsf{pl}_{Z^{n}}(Z_{n+1})\leq t_{n}(\alpha)\}\leq\alpha\quad\text{for all $n$ and all $\alpha\in(0,1)$},$ (2)
where the supremum is over all distributions $\mathsf{P}$ for the exchangeable sequence.
Proof.
See the Appendix. ∎
Theorem 1 implies that the prediction region in (1) is finite-sample (joint) valid in the sense that
$\mathsf{P}^{n+1}\{Y_{n+1}\in C_{\alpha}(X_{n+1};Z^{n})\}\geq 1-\alpha\quad\text{for all $(n,\mathsf{P})$},$ (3)
where $\mathsf{P}^{n+1}$ is the joint distribution for $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n}),(X_{n+1},Y_{n+1})$. Namely, the coverage probability of using the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(x,Z^{n})$ for prediction is no less than $1-\alpha$ for all sample size $n$ and all distribution $\mathsf{P}$. Note that this notion of finite-sample validity of the prediction region is distinct from a more desirable property of conditional validity:
$\mathsf{P}^{n+1}\{Y_{n+1}\in C_{\alpha}(X_{n+1};Z^{n})\mid X_{n+1}=x\}\geq 1-\alpha\quad\text{for all $(n,\mathsf{P})$ and almost all $x$}.$
By the iterated expectation formula, we know conditional validity implies joint validity:
$\mathsf{P}^{n+1}\{Y_{n+1}\in C_{\alpha}(X_{n+1};Z^{n})\}=\mathsf{E}\left[\mathsf{P}^{n+1}\{Y_{n+1}\in C_{\alpha}(X_{n+1};Z^{n})\mid X_{n+1}\}\right],$
where the expectation is taken with respect to the distribution of $X_{n+1}$. Informally, the conditional validity states that the probability of a correct prediction is no less than $1-\alpha$ each time the actuary makes a prediction. In contrast, the joint validity says that the overall rate of correct predictions is at least $1-\alpha$. While conditional validity is more desirable than joint validity, it is practically unachievable. Indeed, Vovk (2012) and Lei and Wasserman (2014) show that it is impossible to achieve conditional validity property with a bounded prediction region $C_{\alpha}(X_{n+1};Z^{n})$ in supervised learning. Hong (2023) establishes a similar result for unsupervised learning. Therefore, the finite-sample joint validity is practically the best we can hope for.
3 Proposed method
In many insurance applications, $Y$ denotes the insurance claims or something alike, and $X$, a $p$-dimensional vector in $\mathbb{R}^{p}$, symbolizes $p$ features. In this paper, we only consider the regression problem, i.e., the case where the response variable $Y$ is continuous; however, see, Section 5 for some remarks for the case of a categorical dependent variable. Therefore, we assume that $Y$ is a non-negative random variable supported on the nonnegative real line $\mathbb{R}_{+}=[0,\infty)$. An insurer often needs to set its risk capital to an adequate level to comply with insurance regulations (e.g., European Solvency II 2009) or implement its risk management plan. Therefore, an insurer needs to generate a finite-sample valid prediction interval of the form $[0,b)$, where the upper bound $b=b(X_{n+1};Z^{n})$ is a positive data-dependent quantity. There is a paucity of methods for achieving this in the insurance literature. In particular, if an actuary applies a linear model, then predictions will be subject to model misspecification risk. Even if the model assumption can be fully justified, the selection effect can still adversely affect the prediction accuracy (e.g., Hong et al. 2018a, b). While existing non-parametric regression methods (e.g., those methods based on kernel smoothing, basis expansion, regression splines, etc) might alleviate these issues, they, however, can only generate asymptotically valid prediction intervals for $Y_{n+1}$, let alone the challenge of choosing the tuning parameters of these models. Fortunately, conformal prediction can come to the rescue.
It is clear from Algorithm 1 and Theorem 1 that conformal prediction is a model-free method. In particular, finite-sample validity, given by (2), holds regardless of the non-conformity measure an actuary chooses. However, to apply conformal prediction, an actuary must overcome several challenges. First, the shape of the conformal prediction region $C_{\alpha}(x,Z^{n})$ depends on the choice of the non-conformity measure. In particular, when the response variable $Y\in\mathbb{R}_{+}$, there is no guarantee that the conformal prediction region $C_{\alpha}(x,Z^{n})$ is an interval of the form $[0,b)$. The shape of $C_{\alpha}(x,Z^{n})$ can be a disjoint union of several intervals; see, for instance, Section 2.3 of Vovk et al. (2005) and Lei et al. (2013). If this happens, the resulting conformal region will not be practically useful. In addition, Algorithm 1 cannot be implemented exactly in general because there are infinitely many provisional values $y$ of $Y$ and an actuary can only consider finitely many of them. This means the implementation of Algorithm 1 is usually approximate only, jeopardizing finite-sample validity—one of the key selling points of conformal prediction. Finally, even for an approximate implementation of Algorithm 1, i.e., we only consider a grid of $y$ values, the computation required could be prohibitively expensive.
To address these challenges together, we consider the following non-conformity measure. Let $B=\{z_{1},\ldots,z_{n}\}$ be a bag of observations of size $n$. For $i=1,\ldots,n$, let $z_{i}=(x_{i1},\ldots,x_{ip},y_{i})$ be the $i$-th observations in $B$. That is, $x_{ij}$ denotes the $i$-th observation of the the $j$-th feature. Suppose $z=(x_{1},\ldots,x_{p},y)$ is a provisional value of $Z_{n+1}$. We define $M$ to be
$M(B,z)=y-\left[\frac{1}{n}\sum_{j=1}^{p}x_{j}+\sum_{i=1}^{n}\left(y_{i}-\frac{1}{n}\sum_{j=1}^{p}x_{ij}\right)\right].$ (4)
In (4), $\sum_{i=1}^{n}(y_{i}-\sum_{j=1}^{p}x_{ij}/n)$ is the only term that depends on the bag $B$. Informally, it measures the aggregate difference between the response $Y$ and the sum of the features per data point. However, it is worth underlining that we are not proposing a model like $Y=X_{1}/n+\ldots+X_{p}/n+\varepsilon$, where $\varepsilon$ is some random error term. Conformal prediction is a model-free method; hence, there is no model to consider. $M$ in (4) is just a non-conformity measure we propose to address the above challenges. This non-conformity measure possesses several desirable properties: (i) the resulting prediction region is an interval of form $[0,b)$, (ii) there is a closed-form formula for the resulting prediction interval, making implementation simple (iii) the implementation is exact, guaranteeing finite-sample validity, and (iv) it allows us to create a tuning-parameter-free and model-free non-parametric distribution function for $Y_{n+1}$. Properties (i)—(iii) immediately follow from the following theorem.
Theorem 2.
Suppose the non-conformity measure is given by (4). For $1\leq i\leq n$, let $W_{i}=Y_{i}+\sum_{i=1}^{p}(X_{(n+1)j}-X_{ij})/n$. Then for any $10<\alpha<$ 1, the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n+1};Z^{n})$ for $Y_{n+1}$ equals $(0,W_{(k)})$, where $k=\min\{n,\lfloor(n+1)(1-\alpha)+1\rfloor\}$.
Proof.
See the Appendix. ∎
Regarding Property (iv), consider the following function $\mathbb{G}_{n}$:
$\mathbb{G}_{n}(y)=\frac{\sum_{i=1}^{n+1}1_{\{\mu_{i}\leq\mu_{n+1}\}}(y)}{n+1},\ y\in\mathbb{R}_{+}.$ (5)
Evidently, $\mathbb{G}_{n}$ does not depend on any model or tuning parameters. The next theorem shows that it is a bona fide distribution function and that it converges to the true distribution of $Y_{n+1}$ as $n\rightarrow\infty$.
Theorem 3.
Suppose $Z_{1},Z_{2},\ldots$ is an iid sequence of observations and $\mathsf{P}$ is the distribution of $Z_{1}$, then the function $\mathbb{G}_{n}$ in (5) is a genuine distribution function and can be written as
$\mathbb{G}_{n}(y)=\frac{\sum_{i=1}^{n+1}1_{\{W_{i}\leq y\}}(y)}{n+1},\ y\in\mathbb{R}_{+},$ (6)
where $W_{i}=Y_{i}+\sum_{i=1}^{p}(X_{(n+1)j}-X_{ij})/n$. Moreover, this conformal predictive distribution function $\mathbb{G}_{n}(y)$ converges to $\mathbb{G}(y)$ both in probability and in distribution for each $y\in\mathbb{R}_{+}$, where $\mathbb{G}$ is the distribution function of $Y_{n+1}$.
Proof.
See the Appendix. ∎
It bears remaking that non-parametric distribution functions based on conformal prediction have been considered by Vovk et al. (2019). However, what Vovk et al. (2019) propose is a general framework for generating a random distribution function, and it is unclear how a random distribution function can be practically useful for predictive inference in insurance. The function $\mathbb{G}_{n}$ in (5) is a deterministic and legitimate distribution function. It can be considered as a specially adapted form of the random conformal distribution function in Vovk (2024), based on our non-conformity measure in (4).
In addition, the convergence of $\mathbb{G}_{n}$ in Theorem 3 is not to be confused with the finite-sample validity of the conformal predictive region $C_{\alpha}(x;Z^{n})$. The latter is guaranteed by Theorem 1, irrespective of the sample size $n$. The former is an asymptotic property of the conformal distribution function $\mathbb{G}_{n}$ which is an estimator of the true distribution function $\mathbb{G}$ of $Y_{n+1}$. Since a distribution function for $Y_{n+1}$ is more informative than a prediction interval for $Y_{n+1}$, $\mathbb{G}_{n}$ might provide an actuary with valuable information when the sample size $n$ is large.
4 Examples
Throughout this section, the following notation will be used:
  1. $\bullet$
    ${\sf Unif}(c,d)$ denotes the uniform distribution supported on $(c,d)$;
  2. $\bullet$
    ${\sf Exp}(\lambda)$ stands for the exponential distribution with rate $\lambda$. That is, the exponential distribution with the probability density function
    $f(x)=\lambda e^{-\lambda x},\quad x>0;$
  3. $\bullet$
    ${\sf LogN}(\mu,\sigma)$ designates the lognormal distribution with parameters $\mu$ and $\sigma$, i.e, the distribution of the random variable $X=e^{\mu+\sigma W}$, where $W$ follows the standard normal distribution;
  4. $\bullet$
    ${\sf Gamma}(a,b)$ represents the gamma distribution with shape parameter $a$ and rate parameter $b$, or equivalently, the gamma distribution whose probability density function is
    $g(x)=\frac{b^{a}}{\Gamma(a)}x^{a-1}e^{-bx},\quad x>0,$
[…]
  5. $\bullet$
    ${\sf Pareto}(\eta,\beta)$ symbolizes the Type II Pareto distribution with the probability density function
    $h(x)=\frac{\eta\beta^{\eta}}{(x+\beta)^{\eta+1}},\quad x>0.$
Example 1
Here we provide a simulation study to demonstrate the performance of the proposed conformal prediction interval. First, assume data is generated from the following $3$-feature linear model:
$Y=2X_{1}+5X_{2}+3X_{3}+\varepsilon,$ (7)
where $\varepsilon$ is a random error term, $X_{1}\sim\exp(5)$, $X_{2}\sim{\sf Unif}(0,1)$, $X_{3}\sim{\sf LogN}(1,2)$, and $X_{1},X_{2},X_{3}$, and $\varepsilon$ are independent. In our simulation, we will consider three distributions for $\varepsilon$: ${\sf LogN}(2,2)$, ${\sf Gamma}(5,0.5)$ , and ${\sf Pareto}(1.5,4)$. Therefore, we have three oracle models. For $\alpha=0.15,0.10$, and $0.05$ and $n=20,50$, and $200$, we generate $N=5,000$ random samples of size $n+1$ from each of these three models. For each sample, we use the values of the response variable in the first $n$ sample points and all the values of the features to construct the $100(1-\alpha)\%$ conformal prediction interval in Theorem 2. Then we use the value of the response variable of the last sample point, i.e, the value of $(n+1)$-st response variable, to test its validity. We estimate the coverage probability of each conformal prediction interval as $K/N$, where $K$ is the number of times it contains the value of the $(n+1)$-st response variable. Tables 13 summarize the results. They show that the proposed conformal prediction interval achieves the nominal coverage probability level across different error distributions (lognormal, gamma, and Pareto), different $\alpha$ levels ($0.85,0.90$, and $0.95$), and different sample sizes ($20,50$, and $100$).
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8622
${\sf LogN}(1,2)$ 0.90 0.9138
0.95 0.9584
0.85 0.8566
${\sf Gamma}(3,0.5)$ 0.90 0.9046
0.95 0.9544
0.85 0.8596
${\sf Pareto}(1.5,4)$ 0.90 0.9088
0.95 0.9560
Table 1: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various values of $\alpha$ and distributions of $\varepsilon$, when the oracle model is given by (7) and the sample size is $n=20$.
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8616
${\sf LogN}(1,2)$ 0.90 0.9018
0.95 0.9596
0.85 0.8570
${\sf Gamma}(3,0.5)$ 0.90 0.9014
0.95 0.9598
0.85 0.8642
${\sf Pareto}(1.5,4)$ 0.90 0.9016
0.95 0.9584
Table 2: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various distributions of $\varepsilon$, when the oracle model is given by (7) and the sample size is $n=50$.
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8516
${\sf LogN}(1,2)$ 0.90 0.9058
0.95 0.9510
0.85 0.8534
${\sf Gamma}(3,0.5)$ 0.90 0.9022
0.95 0.9538
0.85 0.8586
${\sf Pareto}(1.5,4)$ 0.90 0.9076
0.95 0.9538
Table 3: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various values of $\alpha$ and distributions of $\varepsilon$, when the oracle model is given by (7) and the sample size is $n=200$.
Next, we consider the following quadratic model:
$Y=2X_{1}+5X_{2}^{2}+3X_{3}+\varepsilon,$ (8)
where $\varepsilon$ is a random error term, $X_{1}\sim\exp(5)$, $X_{2}\sim{\sf Unif}(0,1)$, $X_{3}\sim{\sf LogN}(1,2)$, and $X_{1},X_{2},X_{3}$, and $\varepsilon$ are independent. We repeat the same simulation as above except that data is generated from (8) instead of (7). The results are summarized in Tables 46. Once again, the proposed conformal prediction interval achieves the nominal coverage probability level in every case scenario.
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8626
${\sf LogN}(1,2)$ 0.90 0.9140
0.95 0.9584
0.85 0.8618
${\sf Gamma}(3,0.5)$ 0.90 0.9070
0.95 0.9514
0.85 0.8064
${\sf Pareto}(1.5,4)$ 0.90 0.9086
0.95 0.9562
Table 4: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various values of $\alpha$ and distributions of $\varepsilon$, when the oracle model is given by (8) and the sample size is $n=20$.
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8616
${\sf LogN}(1,2)$ 0.90 0.9016
0.95 0.9596
0.85 0.8622
${\sf Gamma}(3,0.5)$ 0.90 0.9012
0.95 0.9614
0.85 0.8642
${\sf Pareto}(1.5,4)$ 0.90 0.9018
0.95 0.9582
Table 5: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various values of $\alpha$ and distributions of $\varepsilon$, when the oracle model is given by (8) and the sample size is $n=50$.
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8512
${\sf LogN}(1,2)$ 0.90 0.9058
0.95 0.9510
0.85 0.8538
${\sf Gamma}(3,0.5)$ 0.90 0.9024
0.95 0.9538
0.85 0.8590
${\sf Pareto}(1.5,4)$ 0.90 0.9076
0.95 0.9538
Table 6: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various values of $\alpha$ and distributions of $\varepsilon$, when the oracle model is given by (8) and the sample size is $n=200$.
Finally, we repeat the above simulation with data generated from the following polynomial model:
$Y=2X_{1}+5X_{2}^{2}+3X_{3}^{3}+\varepsilon,$ (9)
where $\varepsilon$ is a random error term, $X_{1}\sim\exp(5)$, $X_{2}\sim{\sf Unif}(0,1)$, $X_{3}\sim{\sf LogN}(1,2)$, and $X_{1},X_{2},X_{3}$, and $\varepsilon$ are independent. Similar to the above two cases, the conformal prediction interval achieves the nominal coverage probability level across different settings.
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8636
${\sf LogN}(1,2)$ 0.90 0.9050
0.95 0.9536
0.85 0.8582
${\sf Gamma}(3,0.5)$ 0.90 0.9068
0.95 0.9540
0.85 0.8576
${\sf Pareto}(1.5,4)$ 0.90 0.9066
0.95 0.9550
Table 7: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various values of $\alpha$ and distributions of $\varepsilon$, when the oracle model is given by (9) and the sample size is $n=20$.
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8716
${\sf LogN}(1,2)$ 0.90 0.9098
0.95 0.9648
0.85 0.8620
${\sf Gamma}(3,0.5)$ 0.90 0.9014
0.95 0.9618
0.85 0.8620
${\sf Pareto}(1.5,4)$ 0.90 0.9052
0.95 0.9588
Table 8: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various values of $\alpha$ and distributions of $\varepsilon$, when the oracle model is given by (9) and the sample size is $n=50$.
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8650
${\sf LogN}(1,2)$ 0.90 0.9108
0.95 0.9562
0.85 0.8558
${\sf Gamma}(3,0.5)$ 0.90 0.9028
0.95 0.9530
0.85 0.8574
${\sf Pareto}(1.5,4)$ 0.90 0.9080
0.95 0.9536
Table 9: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various values of $\alpha$ and distributions of $\varepsilon$, when the oracle model is given by (9) and the sample size is $n=200$.
Example 2
Suppose data is generated from the following model:
$Y=2X_{1}+\varepsilon,$ (10)
where $X_{1}\sim{\sf Exp}(5)$, $\varepsilon\sim{\sf Gamma}(2,2.5)$, and $X_{1}$ and $\varepsilon$ are independent. The closure property of the gamma distribution implies $Y\sim{\sf Gamma}(3,2.5)$. We simulate two samples of different sizes ($n=50$ and $500$) from (10). Then we plot the corresponding conformal distribution function $\mathbb{G}_{n}$ in (5) along with the true distribution function $\mathbb{G}$ of $Y$. Figure 1 shows these plots. In both panels, the dotted line denotes the curve of the true distribution function $\mathbb{G}$ of $Y$ and the solid curve is the curve of the conformal distribution function $\mathbb{G}_{n}$. When the sample size is relatively small, i.e., $n=50$, $\mathbb{G}_{n}$ does not approximate $\mathbb{G}$ well enough. This should not be taken as a shortcoming of the conformal distribution function because no method is expected to approximate the true distribution well for such a small sample size. However, $\mathbb{G}_{n}$ approximates $\mathbb{G}$ very well when the sample size is $n=500$; this is consistent with Theorem 3.
(a) $n=50$
(b) $n=500$
Figure 1: Plots of the true distribution function $\mathbb{G}$ of $Y$ (dotted) and the conformal distribution function $\mathbb{G}_{n}$ (solid) for two different sample sizes in Example 2.
Example 3
Consider the term life insurance data in the example on Page 70 of Frees (2010). This data set contains the survey results of $500$ randomly sampled households with positive incomes. It is available at https://instruction.bus.wisc.edu/jfrees/jfreesbooks. Following Frees (2010), we only consider the subset of $275$ families that purchased term life insurance. The response variable FACE is the amount the insurer will pay in the event of the death of the insured. There are three features: annual INCOME, the number of years of EDUCATION of the surveyed respondent, and the number of household members, NUMHH. Table 10 gives the summary statistics of the data.
Variable Minimum 1st Quartile Median 3rd Quartile Maximum Mean
FACE 800 50,000 150,000 590,000 14,000,000 747,581
INCOME 260 38, 000 65,000 120,000 100,000,00 208,975
EDUCATION 2.00 13.00 16.00 16.50 17.00 14.52
NUMHH 1.00 2.00 3.00 4.00 9.00 2.96
Table 10: Summary statistics of the term life insurance data in Example 3.
Table 11 shows the upper bound of the $100(1-\alpha)\%$ conformal prediction interval in Theorem 2 for three $\alpha$ levels: $0.85,0.90$, and $0.95$. Given the excellent performance of the proposed conformal prediction interval in Example 1, we are inclined to believe that the conformal prediction interval here achieves the nominal coverage probability level.
$1-\alpha$ prediction interval upper bound
0.85 1,229,331
0.90 1,842,664
0.95 3,079,330
Table 11: The upper bounds of the $100(1-\alpha)\%$ conformal prediction interval based on term life insurance data in Example 3.
Figure 2 demonstrates the plot of the corresponding conformal distribution function $\mathbb{G}_{n}$. Since the sample size is $n=275$, $\mathbb{G}_{n}$ is likely close to the true distribution of $Y$ which is unobservable here.
Figure 2: Plots of the conformal distribution function $\mathbb{G}_{n}$ based on the term life insurance data in Example 3.
5 Concluding remarks
We have proposed a new non-conformity measure to create a new prediction interval for the response variable in the regression problem. This prediction interval has multiple merits: (i) it avoids both model misspecification risk and the selection effect, (ii) it has a closed-form formula and hence is easy to compute, and (iii) it guarantees finite-sample validity. Our simulation study shows that this prediction interval performs well across different sample size, different nominal coverage probability levels, and different true distributions. In this paper, we have restricted the domain of the response variable $Y$ to $\mathbb{R}_{+}$. However, it is easy to see that results in this paper extend to the case where $Y\in\mathbb{R}$. Finally, we mention that the method proposed in this paper does not apply to the case of a categorical response variable. Yang and Hong (2024) develop a valid model-free method for predicting a categorical response variable and apply it to tackle the problem of detecting insurance fraud.
Appendix
Proof of Theorem 1
The proof is similar to the proof of its counterpart in the unsupervised learning context (e.g., Hong and Martin 2021). In great detail, $\mu_{1},\ldots,\mu_{n},\mu_{n+1}$ are exchangeable because $Z_{1},Z_{2},\ldots$ are exchangeable. Hence, the rank of $\mu_{n+1}$ follows the discrete uniform distribution on the set $\{1,\ldots,n,n+1\}$. By its definition, the plausibility function $\mathsf{pl}_{Z^{n}}(Z_{n+1})$ is proportional to the rank of $\mu_{n+1}$. Therefore, $\mathsf{pl}_{Z^{n}}(Z_{n+1})$ also follows the discrete uniform distribution on the set $\{1/(n+1),2/(n+1),\ldots,1\}$. For a given $10<\alpha<$ 1, if $(n+1)\alpha$ is an integer, then $\mathsf{P}^{n+1}\{\mathsf{pl}_{n+1}(Z_{n+1})\leq t_{n}(\alpha)\}=\alpha$. Otherwise, we will have $\mathsf{P}^{n+1}\{\mathsf{pl}_{n+1}(Z_{n+1})\leq t_{n}(\alpha)\}\leq t_{n}(\alpha)<\alpha$. Therefore, the finite-sample validity, characterized by (2), always holds.
Proof of Theorem 2
For notational simplicity, we let $S_{i}=\sum_{j=1}^{p}X_{ij}$ for $1\leq i\leq n$. We have
$\mu_{i}$ $=$ $M(Z^{n+1}\backslash Z_{i},Z_{i})=Y_{i}-\left[S_{i}/n+\sum_{k=1,k\neq i}^{n+1}(Y_{k}-S_{k}/n)\right]$
$=$ $Y_{i}-\left[S_{i}/n+\sum_{k=1}^{n}(Y_{k}-S_{k}/n)+(Y_{n+1}-S_{n+1}/n)-(Y_{i}-S_{i}/n)\right]$
$=$ $2(Y_{i}-S_{i}/n)-(Y_{n+1}-S_{n+1}/n)-\sum_{k=1}^{n}(Y_{k}-S_{k}/n),\text{ for $1\leq i\leq n$},$
and
$\mu_{n+1}$ $=$ $M(Z^{n},Z_{n+1})=Y_{n+1}-\left[S_{n+1}/n+\sum_{k=1}^{n}(Y_{k}-S_{k}/n)\right]$
$=$ $(Y_{n+1}-S_{n+1}/n)-\sum_{k=1}^{n}(Y_{k}-S_{k}/n).$
Therefore, $\mu_{i}\geq\mu_{n+1}$ if and only if
$2(Y_{i}-S_{i}/n)-(Y_{n+1}-S_{n+1}/n)\geq(Y_{n+1}-S_{n+1}/n),$
which amounts to saying $\mu_{i}\geq\mu_{n+1}$ if and only if $Y_{n+1}\leq Y_{i}+(_{n+1}-S_{i})/n=W_{i}$ for all $1\leq i\leq n$. Therefore, $Y_{n+1}\leq W_{(k)}$ implies that $Y_{n+1}$ is less than at least $(n+1)-k+1$ $W_{i}$’s. Now the theorem follows from the definitions of $\mathsf{pl}_{z^{n}}(x,y)$ and $C_{\alpha}(x;Z^{n})$.
Proof of Theorem 3
For the first part of the theorem, we need to verify that the function $\mathbb{G}_{n}$ satisfies the following conditions:
  1. (a)
    $0\leq\mathbb{G}_{n}(y)\leq 1$ for all $y\in\mathbb{R}$;
  2. (b)
    $\mathbb{G}_{n}$ is non-decreasing;
  3. (c)
    $\mathbb{G}_{n}$ is right-continuous;
  4. (d)
    $\lim_{y\rightarrow-\infty}\mathbb{G}_{n}(y)=0$ and $\lim_{y\rightarrow\infty}\mathbb{G}_{n}(y)=1$.
The definition of $\mathbb{G}_{n}$ shows that Condition (a) is satisfied. In the proof of Theorem 2, we derived that $\mu_{i}\geq\mu_{n+1}$ if and only if $Y_{n+1}\leq Y_{i}+(S_{n+1}-S_{i})/n$ for all $1\leq i\leq n$. Since $y$ is a provisional value of $Y_{n+1}$, this immediately implies that Conditions (b), (c), and (d) are all satisfied. Moreover, it follows from (5) and the fact $\mu_{i}\leq\mu_{n+1}$ if and only if $Y_{n+1}\geq Y_{i}+(S_{n+1}-S_{i})/n$ that
$\mathbb{G}_{n}(y)=\frac{\sum_{i=1}^{n+1}1_{\{W_{i}\leq y\}}(y)}{n+1},\ y\in\mathbb{R_{+}},$
where $W_{i}=Y_{i}+\sum_{i=1}^{p}(X_{(n+1)j}-X_{ij})/n=Y_{i}+(S_{n+1}-S_{i})/n$.
For the second statement of the theorem, note that it does not follow from a trivial application of the law of large numbers because the term $1_{\{\mu_{i}\leq\mu_{n+1}\}}$ depends on $n$ for each $i=1,\ldots,n$. We will apply the theory of empirical processes to establish this statement. To this end, consider the random functions $\{\hat{f}_{n}\}_{n\geq 1}$ where $\hat{f}_{n}(y)=\hat{f}_{n}(y;Z_{1},\ldots,Z_{n},X_{n+1})=1_{\{Y_{i}+(S_{n+1}-S_{i})/n\leq y\}}$. Then each $\hat{f}_{n}$ takes values in a class of measurable functions =$\mathcal{F}=\{f_{\theta}(y)\}_{\theta\in\Theta}$, where $f_{\theta}(y)=1_{\{y_{i}+\theta\leq y\}}$ and $\Theta\subseteq\mathbb{R}$. Since $(s_{n+1}-s_{i})/n\rightarrow 0$, we may assume $\Theta$ is a compact set in $\mathbb{R}$. Therefore, Example 19.7 of van der Vaart (1998) implies that the function class $\mathcal{F}$ is $\mathsf{P}$-Donsker. Next, we let $f_{0}(y)=1_{\{Y_{i}\leq y\}}$. Then $f_{0}\in L^{2}(\mathsf{P})$. Also,
[…]
$=$ $\mathsf{P}\left(\{Y_{i}+(S_{n+1}-S_{i})/n\leq Y<Y_{i}\}\right)+\mathsf{P}\left(\{Y_{i}\leq Y<Y_{i}+(S_{n+1}-S_{i})/n\}\right),$
which converges to 00 in probability as $n\rightarrow\infty$. It follows from Lemma 19.24 of van der Vaart (1998) that $\mathbb{G}_{n}(y)$ converges to $\mathbb{G}(y)$ in probability and in distribution for all $y\in\mathbb{R}_{+}$. This completes the proof.